Write the expression in rectangular form, x+yi, and in exponential form, rele (1 - 1) 9. The rectangular form of the given ex Just want to know if this is right or wrong.
Answer #1

Given n= 9 2 ce-ing where (a+bh where Urtcu a = b = -1 from (1-1) = √2 o- tant (6) tant (7) = -1/4 Rectangular form for zu u